Pls help me solve this question for what value of n is n+1c3=7(nc2)

Answer by rothauserc(4718)

n+1c3=7(nc2)
n +1c3 = 7nc2
subtract n from both sides of =
1c3 = 7nc2 - n
1c3 = (7c2 -1)n
n = 1c3 / (7c2 -1)